# Consumption Partial Insurance

### v5

@@ Appendix D Some welfare costs of tail income risk
Given our empirical, quasi reduced-form approach to measuring partial insurance, we calculate the welfare costs of income risk outside of a fully specified structural model. Specifically, we draw a sequence of income shocks over the lifetime, we use the different consumption specifications to simulate responses to those shocks, and we use parametric utility to calculate welfare based on those consumption responses. Our focus is on the costs of non-Gaussian relative to (counterfactual) Gaussian income risk, and we are particularly interested in contrasting the implications of linear and nonlinear transmission.
@@
$\zeta_{it}\sim\begin{cases}{\cal N}(\mu_{\zeta,1},\sigma_{\zeta,1})\text{ withprob. }p_{\zeta}\\ {\cal N}(\mu_{\zeta,2},\sigma_{\zeta,2})\text{ with prob. }1-p_{\zeta}\\ \end{cases}\text{ and }~{}v_{it}\sim\begin{cases}{\cal N}(\mu_{v,1},\sigma_{v,1})\text{ with prob. }p_{v}\\ {\cal N}(\mu_{v,2},\sigma_{v,2})\text{ with prob. }1-p_{v}.\\ \end{cases}$
@@
These reflect the linear and quadratic consumption functions (4) and (5) in the paper, with $\mu^{c}_{t}=f(\Delta\mathbf{Z}_{it})$ (deterministic profile) and $\xi_{it}=0$ (mean taste shock). In practice, we let $\mu^{c}_{t}$ equal the average (non-residualized) consumption growth by year, which we estimate in the data. We simulate consumption over a working life of 35 years, starting at age 30.888We also simulate a retirement period of 20 years, starting at age 65, in which absence of income shocks implies $\Delta\ln C_{it}=\mu^{c}_{t}$. We fix the initial condition at age 30 ($\ln C_{it=30}$) to its mean value in the sample. In all the simulations, we restrict $\omega_{it}^{(22)}=0$ as this is never precisely estimated in any cut of the sample. In the baseline, we use the time-invariant homogeneous partial insurance parameters from table 3, but we subsequently relax time-invariance and homogeneity across households.
@@
Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=2$. We simulate consumption for 50,000 households using the time-invariant transmission parameters of table 3. The first row presents the average costs across all households. The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience. #of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
@@
Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=2$. We simulate consumption for 50,000 households using age-specific income parameters (table C.9) and consumption transmission parameters (table 4). The first row presents the average costs across all households. The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience. #of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
@@
Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=2$ and low (panel A) versus high wealth (panel B). Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth over 1999–2019. We simulate consumption for 50,000 households using wealth-specific income and consumption transmission parameters, which, for reasons of brevity, are not reported in the main text. The consumption parameters can be proxied by table 6, which reports wealth-and-education-specific results, i.e., an even finer cut of the sample. The first row presents the average costs across all households in the wealth group. The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience. #of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
@@
All households 100% -0.81 -0.88 0.66
$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.80 6.39 7.82
$\min\{\zeta_{it}\}<-0.70$ 29.11% 6.36 8.65 10.04
$\min\{\zeta_{it}\}<-0.95$ 17.07% 8.54 11.89 13.23
Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=1$ (panel A) and $\gamma=3$ (panel B). We simulate consumption for 50,000 households using the time-invariant consumption transmission parameters of table 3. The first row presents the average costs across all households. The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience. #of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.

### v6

@@ Appendix D Some welfare costs of tail income risk
Given our empirical, quasi reduced-form approach to measuring partial insurance, we calculate the welfare costs of income risk outside of a fully specified structural model. Specifically, we draw a sequence of income shocks over the lifetime, we use the different consumption specifications to simulate responses to those shocks, and we use parametric utility to calculate welfare based on those consumption responses. Our focus is on the costs of non-Gaussian relative to (counterfactual) Gaussian income risk, and we are particularly interested in contrasting the implications of linear versus nonlinear transmission.
@@
[…]
@@
These reflect the linear and quadratic consumption functions (4) and (5) in the paper, with $\mu^{c}_{t}=f(\Delta\mathbf{Z}_{it})$ (deterministic profile) and $\xi_{it}=0$ (mean taste shock). In practice, we let $\mu^{c}_{t}$ equal the average (non-residualized) consumption growth by year, which we estimate in the data. We simulate consumption over a working life of 35 years, starting at age 30.999We also simulate a retirement period of 20 years, starting at age 65, in which absence of income shocks implies $\Delta\ln C_{it}=\mu^{c}_{t}$. We fix the initial condition at age 30 ($\ln C_{it=30}$) to its mean value in the sample. In all the simulations, we restrict $\omega_{it}^{(22)}=0$ as this is never precisely estimated in any cut of the sample. In the baseline, we use the time-invariant homogeneous partial insurance parameters from table 3, but we subsequently relax time-invariance and homogeneity across households.
@@
Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=2$. We simulate consumption for 50,000 households using the time-invariant transmission parameters of table 3. The first row presents the average costs across all households. The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience. #% of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
@@
Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=2$. We simulate consumption for 50,000 households using age-specific income parameters (table C.9) and consumption transmission parameters (table 4). The first row presents the average costs across all households. The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience. #% of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
@@
Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=2$ and low (panel A) versus high wealth (panel B). Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth over 1999–2019. We simulate consumption for 50,000 households using wealth-specific income and consumption transmission parameters, which, for reasons of brevity, are not reported in the main text. The consumption parameters can be proxied by table 6, which reports wealth-and-education-specific results, i.e., an even finer cut of the sample. The first row presents the average costs across all households in the wealth group. The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience. #% of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.
@@
All households 100% -0.76 -0.81 0.66
$\min\{\zeta_{it}\}<-0.50$ 40.31% 4.68 6.32 7.69
$\min\{\zeta_{it}\}<-0.70$ 29.11% 6.22 8.59 9.93
$\min\{\zeta_{it}\}<-0.95$ 17.07% 8.40 11.89 13.17
Notes: The table reports the welfare costs of non-Gaussian relative to Gaussian income risk, for $\gamma=1$ (panel A) and $\gamma=3$ (panel B). We simulate consumption for 50,000 households using the time-invariant consumption transmission parameters of table 3. The first row presents the average costs across all households. The next rows present the average costs over subsets of households defined by the magnitude of the worst permanent shock they experience. #% of households who face a shock at least as bad, drawn from the normal mixture, at least once in their lifetime.